OpenStudy (anonymous):

Write the equation of a line given the following info. (-4, -5) and (-6, 12).

OpenStudy (marigirl):

what we can do first is find the gradient using y2-y1/x2-x1 m=(12+5)/(-6+4) m=(17/-2) m=-17/2 or

OpenStudy (marigirl):

m=-8.5

OpenStudy (marigirl):

then we can use y-y1=m(x-x1) to formulate an linear equation where x1 and y1 are a coordinate. so lets choose (-4, -5) y+5=-8.5(x+4) y+5=-8.5x-34 y=-8.5-39